# Bye bye, local bias: the statistics of the halo field are poorly determined by the local mass density

### v2

@@ Abstract
Bias models relating the dark matter field to the spatial distribution of halos are widely used in current cosmological analyses. Many models predict halos purely from the local Eulerian matter density, yet bias models in perturbation theory require the inclusion of other local properties. We assess the validity of assuming that only the local dark matter density can be used to predict the number density of halos in a model-independent way and in the non-perturbative regime. Utilising $N$-body simulations, we study the properties of the halo counts field after spatial voxels with near-equal dark matter density have been permuted. If local-in-matter-density biasing were valid, the statistical properties of the permuted and un-permuted fields would be indistinguishable since both represent equally fair draws of the stochastic biasing model. For voxels of side length $\sim 4-30\,h^{-1}{\rm\,Mpc}$ and for halos less massive than $\sim 10^{15}\,h^{-1}{\rm\,M_{\sun}}$, we find that the permuted halo field has a scale-dependent bias with greater than 25% more power on scales relevant for current surveys. These bias models remove small-scale power by not modelling correlations between neighbouring voxels, which substantially boosts large-scale power to conserve the field’s total variance. This conclusion is robust to the choice of initial conditions and cosmology. Assuming local-in-matter-density halo biasing cannot, therefore, reproduce the distribution of halos across a large range of scales and halo masses, no matter how complex the model. One must either allow the biasing to be a function of other quantities and/or remove the assumption that neighbouring voxels are statistically independent.
@@ 1 Introduction
We present a model-independent test of the assumption of LIMD halo biasing. Our test is independent of the functional form of the biasing model, but simply tests whether halo biasing can be a stochastic function of only the local matter density. We find that LIMD halo biasing cannot reproduce the distribution of halos less massive than $\sim 10^{15}\,h^{-1}\,{\rm M_{\sun}}$ when applied to gridded dark matter fields with voxels of size less than $\sim 30\,h^{-1}{\rm\,Mpc}$. Our permutation test naturally enforces conservation of all 1-point moments and count statistics, while demonstrating the insufficiency of the LIMD hyopthesis to model halo clustering. We demonstrate that this conclusion is robust across initial conditions and cosmology by utilising the quijote suite of $N$-body simulations. We conclude that no LIMD biasing model, however complex, should be used if one wishes to accurately predict the spatial distribution of halos given a matter density field.
@@ 2.1 Assessing the assumption of local biasing
Note that we will not make any assumptions about the type of distribution (e.g. Poisson) nor the functional form of its parameters (e.g. the mean for a Poisson distribution) other than we expect that this function should be a continuous function of the local dark matter density.
@@ 3.1 LIMD halo biasing cannot reproduce the halo distribution
Figure 2: Fractional error on the power spectrum after permuting equal-density voxels as a function of halo mass. Significant discrepancies are seen for a wide range of scales for halo masses below $10^{15}\,h^{-1}{\rm\,M}_{\sun}$, hence LIMD halo biasing is invalid in this regime. The points and error bars give the mean and standard deviation of the discrepancy across 100 permutations, respectively.
In Fig. 2 we investigate how this conclusion changes as a function of halo mass. One observes that the fractional difference remains relatively constant until $\sim 10^{15}\,h^{-1}{\rm\,M}_{\sun}$, indicating that LIMD halo biasing is invalid for all masses below this point. For the largest halo masses, we find that the power spectrum of the permuted field is consistent with the truth. The number density of these objects is low – there are only 381 halos above $10^{15}\,h^{-1}{\rm\,M}_{\sun}$ in the fiducial simulation – and these can only exist at the extreme peaks of the density field. Hence, given that these objects cannot exists in low density environments, it is perhaps unsurprising that their number density can be well-predicted by just the local density field. We therefore conclude that LIMD halo biasing is a reasonable assumption for halos above $\sim 10^{15}\,h^{-1}{\rm\,M}_{\sun}$, but, for the majority of objects in the simulation, this framework cannot be used.
@@
[…]
@@ 3.2 Sensitivity to binning scheme
We have verified that our results are unchanged if we consider equally spaced bins in both density and the logarithm of density, instead of enforcing each bin to have approximately an equal number of voxels. The transient behaviour disappears at $n_{\rm bin}\approx 10$ when using logarithmic bins and at $n_{\rm bin}\approx 300$ for the linear binning scheme. Above these values, we again find that the bias is insensitive to the choice of $n_{\rm bin}$.
Of course, if one as many bins as number of voxels ($n_{\rm bin}=N^{3})$, then (assuming no voxels have identically equal densities) there would be only one voxel per bin and thus the permuted halo field is equal to the unpermuted field. For $N=128$ this corresponds to $n_{\rm bin}\approx 10^{6.3}$. Since our conclusion holds at $n_{\rm bin}=10^{5}$, even with an average of just 21 voxels per bin we cannot reproduce the halo distribution. A LIMD halo biasing model which varies that rapidly with density seems implausible and highly fine-tuned. We have verified that using $n_{\rm bin}=100$ for a mock halo counts field generated from a Poisson distribution with a mean which depends on density according to a power law (with parameters optimised to fit the quijote simulations using a Poisson loss function in each voxel) yields unbiased $P_{\rm counts}(k)$ upon shuffling, again indicating that our fiducial choice of $n_{\rm bin}=10^{3}$ is sufficient.
@@ 4 Discussion and conclusion
In this letter we have shown that the distribution of halos cannot be modelled according to LIMD biasing: the assumption that the number of halos in a given voxel is drawn from a probability distribution which depends solely on the local dark matter density and where each voxel is statistically independent. At least one of these assumptions must be broken to accurately model the distribution of halos in the Universe when averaging over scales of $\sim 4-30\,h^{-1}{\rm\,Mpc}$.
@@
We conclude that number of halos in a given region of space is a stochastic function of the local dark matter density field, but it cannot depend solely on this quantity. Models which rely on this assumption, no matter how complex, cannot reproduce the distribution of halos and thus must be superseded by more accurate models. This letter not only disproves the LIMD assumption of halo biasing for scales $\lesssim 30\,h^{-1}{\rm\,Mpc}$, but provides a framework to aid the design of future approaches by testing their assumptions in a model-independent manner.
We thank Andrés Balaguera, Yan-Chuan Cai, Simon Ding, Ludvig Doeser, Rozo Eduardo, Pedro Ferreira, Joachim Harnois-Deraps, Jens Jasche, Guilhem Lavaux, Titouan Lazeyras, Stuart McAlpine, Minh Nguyen, Francisco Villaescusa-Navarro and the Implicit Likelihood Inference, Accelerated Forward Modelling and Robustness working groups of the Simons Collaboration on “Learning the Universe” for useful comments and suggestions. We particularly thank Neal Dalal and Cora Uhlemann for independently checking our analysis after it appeared as a pre-print, and prompting us to consider the sensitivity of our results to the mass assignment scheme. DJB and MH are supported by the Simons Collaboration on “Learning the Universe.” BDW acknowleges the DIM-ORIGINES-2023, Infinity Next grant. The Flatiron Institute is supported by the Simons Foundation. This work was done within the Aquila Consortium (https://www.aquila-consortium.org/). For the purposes of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.

### v3

# Bye-bye, Local-in-matter-density Bias: The Statistics of the Halo Field Are Poorly Determined by the Local Mass Density

@@ Abstract
Bias models relating the dark matter field to the spatial distribution of halos are widely used in current cosmological analyses. Many models predict halos purely from the local Eulerian matter density, yet bias models in perturbation theory require other local properties. We assess the validity of assuming that only the local dark matter density can be used to predict the number density of halos in a model-independent way and in the non-perturbative regime. Utilising $N$-body simulations, we study the properties of the halo counts field after spatial voxels with near-equal dark matter density have been permuted. If local-in-matter-density biasing were valid, the statistical properties of the permuted and un-permuted fields would be indistinguishable since both represent equally fair draws of the stochastic biasing model. If the Lagrangian radius is greater than approximately half the voxel size and for halos less massive than $\sim 10^{15}\,h^{-1}{\rm\,M_{\sun}}$, we find the permuted halo field has a scale-dependent bias with greater than 25% more power on scales relevant for current surveys. These bias models remove small-scale power by not modelling correlations between neighbouring voxels, which substantially boosts large-scale power to conserve the field’s total variance. This conclusion is robust to the choice of initial conditions and cosmology. Assuming local-in-matter-density halo biasing cannot, therefore, reproduce the distribution of halos across a large range of scales and halo masses, no matter how complex the model. One must either allow the biasing to be a function of other quantities and/or remove the assumption that neighbouring voxels are statistically independent.
@@ 1 Introduction
We present a model-independent test of the assumption of LIMD halo biasing. Our test is independent of the functional form of the biasing model, but simply tests whether halo biasing can be a stochastic function of only the local matter density. We find that LIMD halo biasing cannot reproduce the distribution of halos less massive than $\sim 10^{15}\,h^{-1}\,{\rm M_{\sun}}$ when applied to gridded dark matter fields, when the Lagrangian radius (Eq. 2) of the halo is larger than approximately half the voxel size. Our permutation test naturally enforces conservation of all 1-point moments and count statistics, while demonstrating the insufficiency of the LIMD hypothesis to model halo clustering. We demonstrate that this conclusion is robust across initial conditions and cosmology by utilising the quijote suite of $N$-body simulations. We conclude that no LIMD biasing model, however complex, should be used if one wishes to accurately predict the spatial distribution of halos given a matter density field.
@@ 2.1 Assessing the assumption of local biasing
Note that we will not make any assumptions about the type of distribution (e.g. Poisson) nor the functional form of its parameters (e.g. the mean for a Poisson distribution) other than we expect that this function should be a continuous function of the local dark matter density. The statement about statistical independence is made within the context of conditional probability. Specifically, while the underlying quantities that influence the distribution in each voxel are indeed correlated among voxels (inherited from the correlation due to large-scale structure), the stochastic quantities drawn from this distribution at each voxel are treated as independent random variables. If the particular draw from the conditional distribution depended on other voxels in any other way than through this inherent large-scale structure correlation, then one would need information beyond the local density field to reconstruct the halo counts field, and thus this would not be a model which is not local in matter density.
@@ 3.1 LIMD halo biasing cannot reproduce the halo distribution
Figure 2: Fractional error on the power spectrum after permuting equal-density voxels as a function of halo mass and voxel size, $\Delta x$. Significant discrepancies are seen for a wide range of scales for halo masses below $10^{15}\,h^{-1}{\rm\,M}_{\sun}$ and when $\Delta x\lesssim 30\,h^{-1}{\rm\,Mpc}$, hence LIMD halo biasing is invalid in this regime. The points and error bars give the mean and standard deviation of the discrepancy across 100 permutations, respectively.
In Fig. 2 we investigate how this conclusion changes as a function of halo mass and voxel size. For the smaller voxel sizes ($\Delta x\lesssim 30\,h^{-1}{\rm\,Mpc}$), one observes that the fractional difference is non-zero until $\sim 10^{15}\,h^{-1}{\rm\,M}_{\sun}$, indicating that LIMD halo biasing is invalid for all masses below this point. For the largest halo masses and voxels, we find that the power spectrum of the permuted field is consistent with the truth. The number density of these objects is low – there are only 381 halos above $10^{15}\,h^{-1}{\rm\,M}_{\sun}$ in the fiducial simulation – and these can only exist at the extreme peaks of the density field. Hence, given that these objects cannot exists in low density environments, it is perhaps unsurprising that their number density can be well-predicted by just the local density field. We therefore conclude that LIMD halo biasing is a reasonable assumption for halos above $\sim 10^{15}\,h^{-1}{\rm\,M}_{\sun}$, but, for the majority of objects in the simulation, this framework cannot be used when the voxel size is below $\sim 30\,h^{-1}{\rm\,Mpc}$.
@@
In Fig. 3 we investigate how the error on $P_{\rm counts}(k)$ varies as a function of grid size for our fiducal mass bin. One observes that the level of disagreement is relatively unchanged on the largest scales as one changes the voxel size from $\sim 4\,h^{-1}{\rm\,Mpc}$ to $\sim 30h^{-1}{\rm\,Mpc}$, indicating that LIMD biasing is inappropriate for density fields smoothed on these scales. As one further increases the voxel size to $\sim 60h^{-1}{\rm\,Mpc}$, the bias starts to diminish, with a bias of $\sim 1.4\sigma$ at $k=0.03\,h\,{\rm Mpc^{-1}}$. In this regime the number of voxels is similar to the number of density bins (as indicated by the black line in Fig. 3), but we have verified that similar behaviour is seen if we reduce $n_{\rm bin}$ to 100. However, even on scales $\sim 0.01\,h{\rm\,Mpc}^{-1}$, we see a mean fractional error of 25% for a voxel size of $\sim 62.5\,h^{-1}{\rm\,Mpc}$. Thus, the LIMD halo biasing is invalid for smoothed density fields with a smoothing scale smaller than $\sim 30h^{-1}{\rm\,Mpc}$, although it may be an acceptable approximation on larger scales.
Figure 4: The significance of the deviation (Eq. 1) between the power spectra of the true and permuted fields as a function of halo mass, for different voxel sizes, $\Delta x$. The dashed line corresponds to $\mathcal{B}=1$. Large values of $\mathcal{B}$ correspond to scenarios where LIMD is a poor approximation.
To further investigate how the error on $P_{\rm counts}(k)$ varies as a function of grid size and mass, we first define a way of summarising the bias across the different wavenumbers. If the fractional error at each $k$ is $\epsilon(k)$ with standard deviation $\sigma(k)$, then we define
$\mathcal{B}^{2}=\frac{\int\left(\frac{\epsilon(k)}{\sigma(k)}\right)^{2}{\rm d}k}{\int{\rm d}k}-1.$ (1)
If the fractional error had a constant mean of $n\sigma(k)$ for all wavenumbers and each wavenumber were independent, then this gives $\langle\mathcal{B}^{2}\rangle=n^{2}$, and thus $\mathcal{B}$ gives a dimensionless estimate of the bias, where $\mathcal{B}=0$ corresponds to being unbiased.
We compute this quantity for the various mass bins and resolutions, and plot the resulting $\mathcal{B}$ as a function of halo mass in Fig. 4. As before, we find that the level of bias is small for $\Delta x\gtrsim 30h^{-1}{\rm\,Mpc}$, with $\mathcal{B}<1$ for all halo masses. For the most massive halos (above $\sim 10^{15}\,h^{-1}{\rm\,M}_{\sun}$), we again see no indication of bias for any grid resolution, presumably since these objects are rare and can only exist in the densest environments. As we move to smaller masses and $\Delta x<30\,h^{-1}{\rm\,Mpc}$, we find that the bias increases, with $\mathcal{B}>8$ for $\Delta x=3.9\,h^{-1}{\rm\,Mpc}$.
Intriguingly, we find that the value of $\mathcal{B}$ does not monotonically increase as one decreases the halo mass, but for $\Delta x=15.6\,h^{-1}{\rm\,Mpc}$, $\mathcal{B}$ reaches a maximum of $\sim 2$ at $\sim 10^{14.2}\,h^{-1}{\rm\,M}_{\sun}$, before decreasing to below one at masses below $\sim 10^{13.4}\,h^{-1}{\rm\,M}_{\sun}$. One can begin to see the same effect for $\Delta x=7.8\,h^{-1}{\rm\,Mpc}$, with the maximum at $\sim 10^{13.5}\,h^{-1}{\rm\,M}_{\sun}$. These maxima occur when the voxel size is approximately twice the Lagrangian radius of the halo, defined such that the halo mass, $M_{\rm h}$, obeys
$M_{\rm h}\equiv\frac{4\pi}{3}\rho_{\rm m}R_{\rm L}^{3},$ (2)
where $\rho_{\rm m}$ is the cosmic mean matter density. Therefore, one would expect to see a similar maximum for $\Delta x=3.9\,h^{-1}{\rm\,Mpc}$ for a halo mass of $\sim 10^{12.4}\,h^{-1}{\rm\,M}_{\sun}$. However, these halos are not resolved in the Quijote simulations, so we cannot test this conjecture here. As one decreases the halo mass below this maximum, the voxel size becomes significantly larger than the Lagrangian radius of the halo, and thus LIMD biasing becomes an increasingly good approximation. If the Lagrangian patch is larger than the voxel size, then information from neighbouring voxels is required to determine the number and positions of halos. Hence, we conclude that there exists an intermediate range of masses for a given voxel size where LIMD biasing breaks down, where the upper limit is due to the rarity of very massive halos, and the lower limit is set by the ability to resolve the Lagrangian radius of individual halos.
@@ 3.2 Sensitivity to binning scheme
We have verified that our results are unchanged if we consider equally spaced bins in both density and the logarithm of density, instead of enforcing each bin to have approximately an equal number of voxels. The transient behaviour disappears at $n_{\rm bin}\approx 10$ when using logarithmic bins and at $n_{\rm bin}\approx 300$ for the linear binning scheme. Above these values, we again find that the bias is insensitive to the choice of $n_{\rm bin}$. We have also verified that our binning is sufficiently fine such that shuffling the density field under this scheme yields a matter power spectrum which is consistent with the unshuffled field.
Of course, if one had as many bins as number of voxels ($n_{\rm bin}=N^{3})$, then (assuming no voxels have identically equal densities) there would be only one voxel per bin and thus the permuted halo field is equal to the unpermuted field. For $N=128$ this corresponds to $n_{\rm bin}\approx 10^{6.3}$. Since our conclusion holds at $n_{\rm bin}=10^{5}$, even with an average of just 21 voxels per bin we cannot reproduce the halo distribution. A LIMD halo biasing model which varies that rapidly with density seems implausible and highly fine-tuned. We have verified that using $n_{\rm bin}=100$ for a mock halo counts field generated from a Poisson distribution with a mean which depends on density according to a power law (with parameters optimised to fit the Quijote simulations using a Poisson loss function in each voxel) yields unbiased $P_{\rm counts}(k)$ upon shuffling, again indicating that our fiducial choice of $n_{\rm bin}=10^{3}$ is sufficient.
@@ 4 Discussion and conclusion
In this letter we have shown that the distribution of halos cannot be modelled according to LIMD biasing: the assumption that the number of halos in a given voxel is drawn from a probability distribution which depends solely on the local dark matter density and where each voxel is statistically independent. At least one of these assumptions must be broken to accurately model the distribution of halos in the Universe when averaging over scales of $\sim 4-30\,h^{-1}{\rm\,Mpc}$. Of course, one would expect that if one went to sufficiently small scales, halos would extend to multiple voxels and thus LIMD biasing must break once the voxels are made sufficiently small. This work quantifies this scale and its relation to the halo’s Lagrangian radius (Eq. 2) using $N$-body simulations.
@@
We conclude that number of halos in a given region of space is a stochastic function of the local dark matter density field, but it cannot depend solely on this quantity. Models which rely on this assumption, no matter how complex, cannot reproduce the distribution of halos and thus must be superseded by more accurate models. This letter not only disproves the LIMD assumption of halo biasing for scales which are not significantly greater than the Lagrangian radius of a halo, but provides a framework to aid the design of future approaches by testing their assumptions in a model-independent manner.
We thank Andrés Balaguera, Yan-Chuan Cai, Neal Dalal, Simon Ding, Ludvig Doeser, Rozo Eduardo, Pedro Ferreira, Joachim Harnois-Deraps, Jens Jasche, Guilhem Lavaux, Titouan Lazeyras, Stuart McAlpine, Minh Nguyen, Cora Uhlemann, Francisco Villaescusa-Navarro and the Implicit Likelihood Inference, Accelerated Forward Modelling and Robustness working groups of the Simons Collaboration on “Learning the Universe” for useful comments and suggestions. DJB and MH are supported by the Simons Collaboration on “Learning the Universe.” BDW acknowleges the DIM-ORIGINES-2023, Infinity Next grant. The Flatiron Institute is supported by the Simons Foundation. This work was done within the Aquila Consortium (https://www.aquila-consortium.org/). For the purposes of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.